# Propeller effect in action: unveiling quenched accretion in the transient X-ray pulsar 4U 0115+63

### v1

@@ 4 Discussion and conclusion
In system where accretion process occurs through the disk, like it is expected in 4U 0115+63, the red noise arises from the stochastic nature of viscosity within the accretion disk (Lyubarskii, 1997). Random fluctuations in viscosity lead to local variations in mass density, which propagate both inwards and outwards through the disk under the influence of viscous diffusion (Mushtukov et al., 2018, 2019). This process generates variability in the mass accretion rate at the inner disk radius, and the fluctuations are subsequently reflected at the NS surface, resulting in the observed variability of the X-ray flux (Revnivtsev et al., 2009). Viscous diffusion acts to suppress mass accretion rate fluctuations at frequencies higher than the local viscous frequency. However, even in a cold accretion disk (Tsygankov et al., 2017a) characterized by low viscosity initial fluctuations are not entirely damped and can still propagate inward, reaching the inner disk radius and influencing the mass accretion rate onto the NS. In systems accreting from a stellar wind, variability in the mass accretion rate is instead driven by the intrinsic inhomogeneity, or clumpiness, of the wind (Walter & Zurita Heras, 2007).
@@
Timing analysis revealed the absence of low-frequency red noise in the PDS of 4U 0115+63 during quiescence (Obs1, Obs2, and Obs2016). This contrasts sharply with Obs3, obtained near periastron passage, where a prominent red-noise component is detected in the PDS, consistent with the presence of accretion, as also indicated by the spectral analysis. As previously shown by Doroshenko et al. (2020), the lack of low-frequency variability is a characteristic feature of non-accreting systems, such as magnetars and radio pulsars.
@@
The results of the joint spectral analysis of Obs1, Obs2, and Obs2016 suggest that the continuum emission of 4U 0115+63 during the quiescent state is well described by a single blackbody model, rather than the two-hump spectrum typically associated with low-level accretion in XRP. This finding aligns with the timing analysis, which also supports a non-accreting scenario. We therefore conclude that, during these observations, the accreting material did not reach the NS surface due to the action of the centrifugal barrier, i.e., the propeller effect.
@@
At the same time, X-ray pulsations at 0.277 Hz, corresponding to the NS spin frequency, are clearly detected in this regime in all three observations. The pulse profiles exhibit a sinusoidal-like shape accompanied by a weak harmonic component and a high root-mean-square pulsed fraction (), consistent with previous XMM-Newton observations (Rouco Escorial et al., 2017, 2020). This pulsation behavior has been interpreted as evidence of non-uniform thermal emission from the NS surface, likely caused by anisotropic cooling due to the presence of a strong magnetic field in the stellar crust (e.g., Geppert et al., 2006; Geppert & Viganò, 2014). Pérez-Azor´ın et al. (2006) modeled the temperature distribution in the crust of a magnetized NS and found that the resulting X-ray spectrum can be well described by a single blackbody with an emitting area significantly smaller than the full stellar surface, consistent with our results. Depending on the crustal temperature distribution for different model configurations, the predicted pulsed fraction from this cooling scenario is in the range of 15–28% for a NS with a magnetic field of G.
@@
The tentative positive correlation between the blackbody temperature and the X-ray flux over the pulse phase (see the two upper panels in Fig. 6) further supports the interpretation of the emission as originating from a cooling atmosphere rather than from ongoing accretion. In a passively cooling, magnetized NS atmosphere, the temperature typically increases with optical depth (see, e.g., Shibanov et al., 1992) meaning that observers looking closer to the surface normal see deeper, hotter layers, while those viewing at larger angles sample cooler, upper layers. This geometry naturally leads to a positive correlation between the observed flux and blackbody temperature across the pulse phase. In contrast, accretion-heated atmospheres are expected to exhibit an inverse temperature profile, with external heating producing hotter upper layers (e.g., Suleimanov et al., 2018). In such a case, the observed flux could be anti-correlated with the apparent temperature, especially in the presence of significant optical depth effects. Therefore, the observed tentative temperature–flux correlation may serve as an indicator of anisotropic thermal cooling rather than ongoing accretion.
@@
On longer timescales, we find that the observed quiescent flux decreased between 2016 and 2023 (see Tables 1 and 6). Such variability on year-long timescales cannot be explained by deep crustal heating, which evolves only over yr (Brown et al., 1998). Instead, it points to processes operating in the outer crust. Similar conclusions were reached by Rouco Escorial et al. (2017, 2020), who studied the low-luminosity states of 4U 0115+63 after its 2015 and 2017 giant outbursts. In that case, the year-to-year differences in flux between post-outburst states can be naturally interpreted as changes in the amount of heat deposited during the preceding accretion episode.
@@
Different authors, based on the magnetic hydrogen atmosphere models of a cooling NS, have theoretically predicted a spectrum similar to a blackbody (e.g., Zane et al., 2000; Ho & Lai, 2001). Due to the low Here we performed also modeling of such emission using undisturbed magnetized hydrogen model atmospheres (Suleimanov et al., 2009). We compute model atmosphere with parameters keV, , 12 km, surface magnetic field normal to the atmosphere of G. The emergent model spectrum is well described with the diluted blackbody with the flux being, , with a color correction factor and a dilution factor . This implies that the actual size of the emission region is larger than that inferred from a simple blackbody model, .
@@
We also investigated the angular dependence of the emergent model emission and find that radiation emitted tangentially to the surface, i.e. along the atmosphere plane, is less shifted to high energies ( for the zenith angle of 81°). It is interesting that the ratio of the observed blackbody temperatures of the pulse-average flux spectrum and the low-flux spectrum of the pulse profile equals to the ratio of the color correction factors for the flux spectrum and the intensity spectrum emitted at : 0.44 keV / 0.39 keV 1.25/1.11 1.13. This means that the observed decrease of blackbody temperature can be explained by the model where at the pulse phase with the lower temperature we see the bright spot inclined at a large angle to the line of sight. Thus, we conclude that a model of the magnetized NS atmosphere can explain the observed spectrum of the source and its phase dependence. More detailed modeling will be presented in a separate paper.
@@
Future observational and theoretical studies of XRPs across different luminosity states will be valuable for advancing our understanding of emission mechanism in highly magnetized NSs in the quiescence. In particular, measurements of polarization properties could reveal the pulsar orientation (see, e.g., Poutanen et al., 2024) and help verify the models proposed above.
@@ Acknowledgements.
- Pérez-Azor´ın et al. (2006) Pérez-Azorín, J. F., Miralles, J. A., & Pons, J. A. 2006, A&A, 451, 1009

### v2

@@ 4 Discussion and conclusion
In systems where the accretion process occurs through the disk, as is expected in 4U 0115+63, the red noise arises from the stochastic nature of viscosity within the accretion disk (Lyubarskii, 1997). Random fluctuations in viscosity lead to local variations in mass density, which propagate both inward and outward through the disk under the influence of viscous diffusion (Mushtukov et al., 2018, 2019). This process generates variability in the mass accretion rate at the inner disk radius, and the fluctuations are subsequently reflected at the NS surface, resulting in the observed variability of the X-ray flux (Revnivtsev et al., 2009). Viscous diffusion acts to suppress mass accretion rate fluctuations at frequencies higher than the local viscous frequency. However, even in a cold accretion disk (Tsygankov et al., 2017a) characterized by low viscosity, initial fluctuations are not entirely damped and can still propagate inward, reaching the inner disk radius and influencing the mass accretion rate onto the NS. In systems accreting from a stellar wind, variability in the mass accretion rate is instead driven by the intrinsic inhomogeneity, or clumpiness, of the wind (Walter & Zurita Heras, 2007).
@@
Timing analysis revealed the absence of low-frequency red noise in the PDSs of 4U 0115+63 during quiescence (Obs1, Obs2, and Obs2016). This contrasts sharply with Obs3, obtained near periastron passage, where a prominent red-noise component is detected in the PDS, consistent with the presence of accretion, as also indicated by the spectral analysis. As previously shown by Doroshenko et al. (2020), the lack of low-frequency variability is a characteristic feature of non-accreting systems, such as magnetars and radio pulsars.
@@
The results of the joint spectral analysis of Obs1, Obs2, and Obs2016 suggest that the continuum emission of 4U 0115+63 during the quiescent state is well described by a single blackbody model, rather than the two-hump spectrum typically associated with low-level accretion in XRPs. This finding aligns with the timing analysis, which also supports a non-accreting scenario. We therefore conclude that, during these observations, the accreting material did not reach the NS surface due to the action of the centrifugal barrier, i.e., the propeller effect.
@@
At the same time, X-ray pulsations at 0.277 Hz, corresponding to the NS spin frequency, are clearly detected in this regime in all three observations. The pulse profiles exhibit a sinusoidal-like shape accompanied by a weak harmonic component and a high root-mean-square pulsed fraction (), consistent with previous XMM-Newton observations (Rouco Escorial et al., 2017, 2020). This pulsation behavior has been interpreted as evidence of nonuniform thermal emission from the NS surface, likely caused by anisotropic cooling due to the presence of a strong magnetic field in the stellar crust (e.g., Geppert et al., 2006; Geppert & Viganò, 2014). Pérez-Azorín et al. (2006) modeled the temperature distribution in the crust of a magnetized NS and found that the resulting X-ray spectrum can be well described by a single blackbody with an emitting area significantly smaller than the full stellar surface, consistent with our results. Depending on the crustal temperature distribution for different model configurations, the predicted pulsed fraction from this cooling scenario is in the range 15–28% for a NS with G.
@@
The tentative positive correlation between the blackbody temperature and the X-ray flux over the pulse phase (see the two upper panels in Fig. 6) further supports the interpretation of the emission as originating from a cooling atmosphere rather than from ongoing accretion. In a passively cooling, magnetized NS atmosphere, the temperature typically increases with optical depth (see, e.g., Shibanov et al., 1992), meaning that observers looking closer to the surface normal see deeper, hotter layers, while those viewing at larger angles sample cooler, upper layers. This geometry naturally leads to a positive correlation between the observed flux and blackbody temperature across the pulse phase. In contrast, accretion-heated atmospheres are expected to exhibit an inverse temperature profile, with external heating producing hotter upper layers (e.g., Suleimanov et al., 2018). In such a case, the observed flux could be anticorrelated with the apparent temperature, especially in the presence of significant optical depth effects. Therefore, the observed tentative temperature–flux correlation may serve as an indicator of anisotropic thermal cooling rather than ongoing accretion.
@@
On longer timescales, we find that the observed quiescent flux decreased between 2016 and 2023 (see Tables 1 and 6). Such variability on year-long timescales cannot be explained by deep crustal heating, which evolves only over – yr (Brown et al., 1998). Instead, it points to processes operating in the outer crust. Similar conclusions were reached by Rouco Escorial et al. (2017, 2020), who studied the low-luminosity states of 4U 0115+63 after its 2015 and 2017 giant outbursts. In that case, the year-to-year differences in flux between post-outburst states can be naturally interpreted as changes in the amount of heat deposited during the preceding accretion episode.
@@
Different authors, based on the magnetic hydrogen atmosphere models of a cooling NS, have theoretically predicted a spectrum similar to a blackbody (e.g., Zane et al., 2000; Ho & Lai, 2001). We also modeled such emission using undisturbed magnetized hydrogen model atmospheres (Suleimanov et al., 2009). We computed a model atmosphere with parameters keV, , 12 km, and a surface -field normal to the atmosphere of G. The emergent model spectrum is described well by a diluted blackbody with the flux , a color correction factor , and a dilution factor . This implies that the actual size of the emission region is larger than that inferred from a simple blackbody model, .
@@
We also investigated the angular dependence of the emergent model emission and find that radiation emitted tangentially to the surface, i.e., along the atmosphere plane, is less shifted to high energies ( for the zenith angle, , of 81°). It is interesting that the ratio of the observed blackbody temperatures of the pulse-average flux spectrum and the low-flux spectrum of the pulse profile is approximately equal to the ratio of the color correction factors for the flux spectrum and the intensity spectrum emitted at : .
This means that the observed decrease in blackbody temperature can be explained by the model in which, at the pulse phase with the lower temperature, we see the bright spot inclined at a large angle to the line of sight. Thus, we conclude that a model of the magnetized NS atmosphere can explain the observed spectrum of the source and its phase dependence. More detailed modeling will be presented in a separate paper.
@@
Future observational and theoretical studies of XRPs across different luminosity states will be valuable for advancing our understanding of emission mechanisms in highly magnetized NSs in quiescence. In particular, measurements of polarization properties could reveal the pulsar orientation (see, e.g., Poutanen et al., 2024) and help verify the models proposed above.
@@ Acknowledgements.
- Pérez-Azorín et al. (2006) Pérez-Azorín, J. F., Miralles, J. A., & Pons, J. A. 2006, A&A, 451, 1009